# Automatic Input Feature Relevance via Spectral Neural Networks

### v2

@@ Abstract
In machine learning practice it is often useful to identify relevant input features, so as to obtain compact dataset for more efficient numerical handling. On the other hand, by isolating key input elements, ranked according their respective degree of relevance, can help to elaborate on the process of decision making. Here, we propose a novel method to estimate the relative importance of the input components for a Deep Neural Network. This is achieved by leveraging on a spectral re-parametrization of the optimization process. Eigenvalues associated to input nodes provide in fact a robust proxy to gauge the relevance of the supplied entry features. Notably, the spectral features ranking is performed automatically, as a byproduct of the network training, with no additional processing to be carried out. The technique is successfully challenged against both synthetic and real data.
@@ 1 Introduction
Deep Neural networks (DNNs) [1, 2, 3, 4, 5] are among the most popular machine learning tools, and, as such, routinely employed in a vast plethora of applications [6, 7, 8, 9]. As for other machine learning techniques, and when employed for e.g. classification and/or regression tasks, DNNs can automatically extract valuable information from the supplied input vector. This amounts to isolate crucial components to implement unmanned decision making strategies. In many cases of interest, input data are multidimensional. In general, it is hence difficult to selectively identify the subset of attributes which prove crucial for the analysis. To identify key input features could be indeed relevant for (i) devising apt solutions for data compression and compactification; (ii) providing critical insights into the processes that led to the produced output.
A first solution is to operate an a priori feature selection, with the aim to isolate beforehand a limited subset of supposedly relevant input elements that should undergo subsequent scrutiny [10]. This is certainly useful to reduce the size of the input to be eventually processed. On the other hand, by clipping part of the dataset, a non zero risk always exists that could yield an unpredictable loss of performance for the trained network. This is as opposed to other viable approaches designed so as to preserve the total amount of information across training.
Among the methods proposed to shed light on the pathways that form the basis of decision of a trained model, worth mentioning is the so called Layer-Wise Relevance Propagation method [11, 12]. This latter implements a backprogation, from the output to the input space, to reach the relevant features that are provided at the entry layer. The method has been successfully challenged against distinct domains of application [13, 14]. Reference should also be made to SHAP [15] (SHapley Additive exPlanations) a game theoretic approach based on the classic Shapley values, which connects optimal credit allocation to local output explanation. Other examples include - but are not limited to - the techniques identified via the compact acronyms IG, DeepLIFT, LIME [16, 17, 18, 19, 20, 21]. To the best of our knowledge, the vast majority of existing methods are however input-dependent. As such, they require a certain number of operations to be carried out after training, for each examined input, so as to assess the relevance of the features for any given supplied item.
@@
Here, we put forward a different method which yields a global metric to assess the relevance of input features to a general feedforward neural network. The procedure applies to the whole dataset and not just to individually examined items. Furthermore, the evaluation of the features is obtained as a byproduct of the training with no additional processing to be carried out. At the end of the training a ranking of the supplied feature is returned, as determined by a set of scalar parameters (one for each distinct feature provided as an input). As mentioned, and unlike the aforementioned methods, the proposed metric is not input dependent. At variance, it provides a global insight on how the model oversees the input space, upon learning. Also, and as opposed to standard techniques for input feature selection, the analysis is not conducted a priori on the data to be eventually scanned.
@@
To reach this goal, we will make use of a recently introduced scheme for handling the optimization of feedforward networks, known as spectral parametrization [22, 23, 24, 25, 26, 27]. The mathematical foundation of the spectral methodology lies within the theory of networks. From this angle, feedforward neural networks can be pictured via a dedicated spectral decomposition that materializes, in direct space, in a collection of adjacency matrices for the linear mapping of the signal across adjacent layers. As we shall argue, the optimized eigenvalues, as made accessible at the end of the training, will return an automatic and self-consistent ranking of the individual input features.
@@
The paper is organized as follows. In the next Section we will review the mathematical basis of the spectral approach to neural network training and introduce the propose ranking strategy. Then, we will turn to discussing a first application to independent and correlated Gaussian data. These are mock data, with different level of complexity, to be regressed by the spectral neural network. In the subsequent Section the proposed method is challenged against the celebrated MNIST database, operated in different modalities. As we will show, the spectral based feature detection method is able to isolate a subset of relevant pixels that prove central for image classification. Then, we turn to challenging the method against an additional dataset, made of stellar spectra. The goal here is to obtain information on relevant frequency bands for the assessment of stellar metallicity. Finally, we sum up and draw our conclusions.
2 On the spectral parametrization
@@
Let us begin by reviewing the basic of the spectral method [22]. We denote with $l-1$ and $l$ two consecutive layers of a feedforward neural network. Assume the layers to be respectively composed by $N$ and $M$ nodes. The vector $\textbf{x}^{l-1}=\left(x^{l-1}_{1},\ldots,x^{l-1}_{N}\right)$ (resp. $\textbf{x}^{l}=\left(x^{l}_{1},\ldots,x^{l}_{M}\right)$) stores information on the activity of the nodes belonging to layer $l-1$ (resp. $l$). The linear transfer between the two adjacent layers is operated by a rectangular $M\times N$ matrix $W$, whose elements read $w_{ij}$. Hence, $\textbf{x}^{l}=f(W\textbf{x}^{l-1})$, where $f$ is a non-linear filter. Under the spectral paradigm, the above operation can be recast in an equivalent form that assumes dealing with the bipartite direct network made of the $N+M$ nodes involved in the transfer. The state vector at the departure layer can be written as $\textbf{z}^{l-1}=\textbf{x}^{l-1}=\left(x^{l-1}_{1},x^{l-1}_{2},\ldots x^{l-1}_{N},0,0,\ldots 0\right)$ where we made explicit that the non trivial activity is solely localized on the nodes of the first layer. Then, $\textbf{z}^{l}=f(A\textbf{z}^{l-1})$, where the last $M$ elements of vector $\textbf{z}^{l}$ refer to the activity on the destination layer; $A$ is a $(M+N)\times(M+N)$ matrix with a $M\times N$ sub-diagonal block that coincides matrix $W$, as illustrated in the Methods section. It should be noted that the diagonal entries of matrix $A$ do not reflect on the produced activity on layer $l$, see Figure 1. One can hence invoke a spectral decomposition of the square transfer matrix, as $A=\Phi\Lambda\Phi^{-1}$ where (i) $\Phi$, the matrix of the eigenvectors, is lower-block triangular with ones on the diagonal, (ii) $\Lambda$ identifies the diagonal matrix formed with the eigenvalues ($\lambda_{i}$, $i=1,\ldots N+M$) of $A$. Given the specific form of matrix $\Phi$, it can be shown [26, 23] that $\Phi^{-1}=2\mathbb{I}-\Phi$. In practical terms, the elements of $A$, namely those involved in the self-consistent definition of the relevant transfer matrix $W$, can be written as a function of the eigenvalues $\lambda_{i}$ and the non trivial entries $\psi_{ij}$ of the eigenvector matrix $\Phi$. An explicit calculation can be performed that yields $w_{i\rightarrow j}=w_{ji}=(\lambda_{i}-\lambda_{j})\psi_{ji}$, where $i=1,...,N$ refer to nodes located on layer $l-1$ and $j=1,...,M$ to those assigned to layer $l$. Further technical details are provided in the Methods section. A direct correspondence can be hence drawn between individual eigenvalues and nodes within the sampled collection. Training a feedforward network under the spectral standpoint amounts to minimize the loss (a suitable metric that gauges the distance between produced and expected output at the exit layer) with respect to the eigenvalues and eigenvectors. In light of the existing correspondence, the magnitude of the optimized eigenvalues can be used as a reliable marker of the nodes’ relevance, an idea that has been explored in [26, 23, 27]. An effective trimming strategy was in particular proposed in [26] which enables for the trained network to be considerably shrunk in size by removing nodes flagged as unimportant. By leveraging on the eigenvalues and their associated magnitudes, it was possible in [27] to estimate the hidden dimension of the target objective function. By expanding further on these ideas, we will hereafter propose an effective input feature detection algorithm. To this end, we chose to operate under the simplifying assumption $\lambda_{j}=0$ (i.e. by forcing to zero the eigenvalues associated to destination nodes) yielding:
@@
The spectral rewriting of the inter-nodes weights $w_{ij}$ yielded therefore a factor, the eigenvalue $\lambda_{i}$, which multiplies all the links emanating from the departure node $i$. The larger the magnitude of the optimized $\lambda_{i}$ the more relevant the associated node $i$ for the processing of the information across the network. Building on this observation, we will show that the trained eigenvalues associated to the starting layer $l=1$ can be effectively employed to automatically discriminate between irrelevant and relevant input features. More specifically, we will demonstrate, that eigenvalues (multiplied by the norm of the corresponding eigenvector entry) can be exploited to rank the relative importance of the features otherwise deemed as irrelevant. The method will be tested against classification problems constructed from mock and real datasets. For all examined cases, a penalty is added to the loss function as in the spirit of [27, 28]. This is a $L_{2}$ regularization of $\lambda_{i}$ which forces eigenvalues associated with irrelevant components to be set to zero (and not just left unaltered to their initial values, possibly drowned in the sea of their relevant homologue). It is worth stressing that in principle it is enough to have just the first layer parameterized as spectral for this analysis to apply. For further details on the training procedure refers to the Methods section.
$l-1$$l$$l-1$$l$$l-1$$l$$l-1$$l$
$=$$A$$\textbf{z}^{l-1}$$\textbf{z}^{l}$$=$$A$$\textbf{z}^{l-1}$$\textbf{z}^{l}$
@@
3 Numerical experiments
The method discussed above for input feature detection is hereafter challenged against different datasets. We will begin in particular by considering a simple multidimensional Gaussian dataset with independent entries. Then, we will turn to considering mock data generated from correlated Gaussian variables. After that, we will move forward to work with the celebrated MNIST dataset and, finally, we will focus on an experimental catalogue of stellar spectra relevant for astrophysical applications
@@ 3.1 Independent Gaussian distributions dataset
The first dataset is structured with 2 target classes and input vectors with 20 independent features. Each data component is sampled from a Gaussian distribution, see Fig. 2. While some components are drawn from distributions which are set to be identical across classes, others display only partial overlap. Components with identical distributions offer no discriminatory information. As such, they prove ineffective for classification purposes. Conversely, the other components are pivotal, the smaller the overlap the larger their relevance in promoting the sought discrimination.
@@
A spectral feedforward neural network, parametrized as stipulated by Eq. (1), has been trained with the aim of assigning input from the above dataset to the correct class of origin. The employed network consists of three hidden layers with, respectively, 100, 100, and 50 neurons. In addition, the network accommodates for an input layer of size 20 and an output layer of size 2. At the end of the training, we analyzed the 20 eigenvalues, associated with the input layer as in the spirit of the above discussion (with an abuse of language we will from here on call eigenvalues the rescaled version of this latter which incorporate the norm of the eigenvectors, as stipulated in the Methods section). As shown in the left panel of Fig. 3, eigenvalues entangled with irrelevant features are all zero, while eigenvalues which can be traced back to relevant input components are non-zero. Hence, the eigenvalues associated to the input layer can efficiently tag features identified as relevant by the network itself to solve the assigned task. The magnitude of non-zero eigenvalues, i.e., those associated with relevant components, is found to correlate with a direct measure of the relative distance between the distributions from which individual components are drawn (see the right panel of Fig. 3). Since the distributions differ only in their mean value, we quantify the above distance as the absolute value of the difference between the two means.

### v3

# Estimating Global Input Relevance and Enforcing Sparse Representations with a Scalable Spectral Neural Network Approach

@@ Abstract
In machine learning practice it is often useful to identify relevant input features. Isolating key input elements, ranked according their respective degree of relevance, can help to elaborate on the process of decision making. Here, we propose a novel method to estimate the relative importance of the input components for a Deep Neural Network. This is achieved by leveraging on a spectral re-parametrization of the optimization process. Eigenvalues associated to input nodes provide in fact a robust proxy to gauge the relevance of the supplied entry features. Notably, the spectral features ranking is performed automatically, as a byproduct of the network training, with no additional processing to be carried out. Moreover, by leveraging on the regularization of the eigenvalues, it is possible to enforce solutions making use of a minimum subset of the input components, increasing the explainability of the model and providing sparse input representations. The technique is compared to the most common methods in the literature and is successfully challenged against both synthetic and real data.
@@ 1 Introduction
Deep Neural networks (DNNs) [bishop2023deep, goodfellow2016deep, lecun2015deep, prince2023understanding, carleo2019machine] are among the most popular machine learning tools, and, as such, routinely employed in a vast plethora of applications [reiser2022graph, chmiela2017machine, chicchi2023frontiers, biancalani2021deep, baldi2014searching]. As for other machine learning techniques, and when employed for e.g. classification and/or regression tasks, DNNs can automatically extract valuable information from the supplied input vector. This amounts to isolate crucial components to implement unmanned decision making strategies. In many cases of interest, input data are multidimensional. In general, it is hence difficult to selectively identify the subset of attributes which prove crucial for the analysis. Moreover, it is also challenging to determine whether the model’s decisions are based on a large portion of the input components or just a minimal subset. To identify key input features could be indeed relevant for (i) devising apt solutions for data compression and compactification; (ii) providing critical insights into the processes that led to the produced output.
Numerous methods have been proposed to shed light on the pathways that form the basis of decision of a trained model [zhang2021survey, dwivedi2023explainable, shrikumar2017learning, swartout1993explanation, montavon2018methods, baehrens2010explain]. Some examples include - but are not limited to - LRP ( Layer-Wise Relevance Propagation) [bach2015pixel, montavon2019layer, bohle2019layer, yang2018explaining], SHAP (SHapley Additive exPlanations)[lundberg2017unified], IG (Integrated Gradients) [sundararajan2017axiomatic] and LIME (Local Interpretable Model-Agnostic Explanations) [lime]. The Layer-Wise Relevance Propagation method backprogates the relevance, a quantity that has multiple definition depending on different rules, from the output to the input space, to reach the relevant features that are provided at the entry layer. SHAP is a game theoretic approach based on the classic Shapley values, which connects optimal credit allocation to local output explanation. IG is a gradient-based approach and, finally, LIME supplies information about features relevance by approximating the model with a local, interpretable model.
To the best of our knowledge, the vast majority of existing methods are however input-dependent. As such, they require a certain number of operations to be carried out after training, for each examined input, so as to assess the relevance of the features for any given supplied item. For most of the aforementioned methods, estimating the global relevance of the input components requires additional computation, typically involving averaging over the entire dataset. More importantly, none of these methods provides a straightforward way to guide the model during training to reach a solution that relies on either a large or a small subset of input components, as they operate on post-training models and play no role during the training phase. This latter property could be especially valuable when deep neural networks are applied to regression tasks in various scientific domains [belfiore2025machine, novakovsky2023obtaining, chicchi2023frontiers, ginolfi2025inferring, camaiani2022model]. In fact, in some research fields, understanding which experimental observations are relevant—and how they are used by the model to solve the task—is sometimes more important than having an opaque model that makes accurate predictions. Identifying a minimal set of input components necessary for solving the task and forcing the model to operate on that sparse representation can help uncover natural relationships among those components.
@@
Here, we put forward a different method which yields a global metric to assess the relevance of input features to a general feedforward neural network. The procedure applies to the whole dataset and not just to individually examined items. Furthermore, the evaluation of the features is obtained as a byproduct of the training with no additional processing to be carried out. At the end of the training a ranking of the supplied feature is returned, as determined by a set of scalar parameters (one for each distinct feature provided as an input). As mentioned, and unlike the aforementioned methods, the proposed metric is not input dependent. At variance, it provides a global insight on how the model oversees the input space, upon learning.
It is important to emphasize that the method we propose is not intended to outperform existing techniques in the literature, some of which remain preferable, especially when the goal is to understand the decision-making process for a specific input. However, the proposed algorithm differs fundamentally from traditional approaches. In fact, it relies on parameters that play a crucial role both during the training phase and the successive process aimed at assessing feature relevance. The integration of these two aspects within a single framework equips the system with the inherent ability to steer the model toward solutions that rely on a reduced number of input components, thereby identifying a minimal subset of features sufficient to efficiently solve the task. As already stated, this capability can be extremely valuable in contexts where explainability is critical (e.g., science, medicine), and where enforcing explainable solutions from the training phase may be the preferred approach.
@@
To reach this goal, we will make use of a recently introduced scheme for handling the optimization of feedforward networks, known as spectral parametrization [giambagli2021machine, chicchi2021training, chicchi2023recurrent, chicchi2024complex, buffoni2022spectral, giambagli2024student]. The mathematical foundation of the spectral methodology lies within the theory of networks. From this angle, feedforward neural networks can be pictured via a dedicated spectral decomposition that materializes, in direct space, in a collection of adjacency matrices for the linear mapping of the signal across adjacent layers. As we shall argue, the optimized eigenvalues, as made accessible at the end of the training, will return an automatic and self-consistent ranking of the individual input features.
@@
The paper is organized as follows. In the next Section we will review the mathematical basis of the spectral approach to neural network training and introduce the propose ranking strategy. Then, we will turn to discussing a first application to independent and correlated Gaussian data. These are mock data, with different level of complexity, to be regressed by the spectral neural network. In the subsequent Section the proposed method is challenged against the celebrated MNIST database, operated in different modalities. As we will show, the spectral based feature detection method is able to isolate a subset of relevant pixels that prove central for image classification. Then, we turn to challenging the method against an additional dataset, made of stellar spectra. The goal here is to obtain information on relevant frequency bands for the assessment of stellar metallicity. A section is then dedicated to compare our method with some of the most commonly used techniques for inferring input feature relevance. As we will show, the spectral scores are in good agreement with those obtained using the other techniques considered. Finally, we sum up and draw our conclusions.
1.1 On the spectral parametrization
@@
Let us begin by reviewing the basic of the spectral method [giambagli2021machine]. We denote with l1l-1 and $l$ two consecutive layers of a feedforward neural network. Assume the layers to be respectively composed by $N$ and $M$ nodes. The vector $\textbf{x}^{l-1}=\left(x^{l-1}_{1},\ldots,x^{l-1}_{N}\right)$ (resp. $\textbf{x}^{l}=\left(x^{l}_{1},\ldots,x^{l}_{M}\right)$) stores information on the activity of the nodes belonging to layer l1l-1 (resp. $l$). The linear transfer between the two adjacent layers is operated by a rectangular $M\times N$ matrix $W$, whose elements read $w_{ij}$. Hence, $\textbf{x}^{l}=f(W\textbf{x}^{l-1})$, where $f$ is a non-linear function. Under the spectral paradigm, the above operation can be recast in an equivalent form that assumes dealing with the bipartite direct network made of the $N+M$ nodes involved in the transfer. The state vector at the departure layer can be written as $\textbf{z}^{l-1}=\textbf{x}^{l-1}=\left(x^{l-1}_{1},x^{l-1}_{2},\ldots x^{l-1}_{N},0,0,\ldots 0\right)$ where we made explicit that the non trivial activity is solely localized on the nodes of the first layer. Then, $\textbf{z}^{l}=f(A\textbf{z}^{l-1})$, where the last $M$ elements of vector $\textbf{z}^{l}$ refer to the activity on the destination layer; $A$ is a $(M+N)\times(M+N)$ matrix with a $M\times N$ sub-diagonal block that coincides matrix $W$, as illustrated in the Methods section. It should be noted that the diagonal entries of matrix $A$ do not reflect on the produced activity on layer $l$, see Figure 1. One can hence invoke a spectral decomposition of the square transfer matrix, as $A=\Phi\Lambda\Phi^{-1}$ where (i) $\Phi$, the matrix of the eigenvectors, is lower-block triangular with ones on the diagonal, (ii) $\Lambda$ identifies the diagonal matrix formed with the eigenvalues ($\lambda_{i}$, $i=1,\ldots N+M$) of $A$. Given the specific form of matrix $\Phi$, it can be shown [buffoni2022spectral, chicchi2021training] that $\Phi^{-1}=2\mathbb{I}-\Phi$. In practical terms, the elements of $A$, namely those involved in the self-consistent definition of the relevant transfer matrix $W$, can be written as a function of the eigenvalues $\lambda_{i}$ and the non trivial entries $\psi_{ij}$ of the eigenvector matrix $\Phi$. An explicit calculation can be performed that yields $w_{i\rightarrow j}=w_{ji}=(\lambda_{i}-\lambda_{j})\psi_{ji}$, where i=1,,Ni=1,...,N refers to nodes located on layer l1l-1 and j=1,,Mj=1,...,M to those assigned to layer $l$. Further technical details are provided in the Methods section. A direct correspondence can be hence drawn between individual eigenvalues and nodes within the sampled collection. Training a feedforward network under the spectral standpoint amounts to minimize the loss (a suitable metric that gauges the distance between produced and expected output at the exit layer) with respect to the eigenvalues and eigenvectors. In light of the existing correspondence, the magnitude of the optimized eigenvalues can be used as a reliable marker of the nodes’ relevance, an idea that has been explored in [buffoni2022spectral, chicchi2021training, giambagli2024student]. An effective trimming strategy was in particular proposed in [buffoni2022spectral] which enables for the trained network to be considerably shrunk in size by removing nodes flagged as unimportant. By leveraging on the eigenvalues and their associated magnitudes, it was possible in [giambagli2024student] to estimate the hidden dimension of the target objective function. By expanding further on these ideas, we will hereafter propose an effective input feature detection algorithm. To this end, we chose to operate under the simplifying assumption $\lambda_{j}=0$ (i.e. by forcing to zero the eigenvalues associated to destination nodes) yielding:
@@
The spectral rewriting of the inter-nodes weights $w_{ij}$ yielded therefore a factor, the eigenvalue $\lambda_{i}$, which multiplies all the links emanating from the departure node $i$. The larger the magnitude of the optimized $\lambda_{i}$ the more relevant the associated node $i$ for the processing of the information across the network. Building on this observation, we will show that the trained eigenvalues associated to the starting layer $l=1$ can be effectively employed to automatically discriminate between irrelevant and relevant input features. More specifically, we will demonstrate, that eigenvalues (multiplied by the norm of the corresponding eigenvector entry) can be exploited to rank the relative importance of the features otherwise deemed as irrelevant. The method will be tested against classification problems constructed from mock and real datasets. For all examined cases, a penalty is added to the loss function as in the spirit of [giambagli2024student, giambagli2024jstat]. This is a $L_{2}$ or a $L_{1}$ regularization of $\lambda_{i}$ which forces eigenvalues associated with irrelevant components to be set to zero (and not just left unaltered to their initial values, possibly drowned in the sea of their relevant homologue). It is worth stressing that in principle it is enough to have just the first layer parameterized as spectral for this analysis to apply. For further details on the training procedure refers to the Methods section.
$l1l-1lll1l-1ll$
[…]
@@
2 Results
The method discussed above for input feature detection is hereafter challenged against different datasets. We will begin in particular by considering a simple multidimensional Gaussian dataset with independent entries. Then, we will turn to considering mock data generated from correlated Gaussian variables. After that, we will move forward to work with the celebrated MNIST dataset and, finally, we will focus on an experimental catalogue of stellar spectra relevant for astrophysical applications.
@@ 2.1 Independent Gaussian distributions dataset
The first dataset is structured with 2 target classes and input vectors with 20 independent features. Each data component is sampled from a Gaussian distribution, see Fig. 2. While some components are drawn from distributions which are set to be identical across classes, others display only partial overlap. Components with identical distributions offer no discriminatory information. As such, they prove ineffective for classification purposes. Conversely, the other components are pivotal, the smaller the overlap the larger their relevance in promoting the sought discrimination. For this test, we generated $10^{5}$ examples.
@@
A spectral feedforward neural network, parametrized as stipulated by Eq. (1), has been trained with the aim of assigning input from the above dataset to the correct class of origin. The employed network consists of three hidden layers with, respectively, 100, 100, and 50 neurons. In addition, the network accommodates for an input layer of size 20 and an output layer of size 2. At the end of the training, we analyzed the 20 eigenvalues, associated with the input layer as in the spirit of the above discussion (with an abuse of language we will from here on call eigenvalues the rescaled version of this latter which incorporate the norm of the eigenvectors, as stipulated in the Methods section). As shown in the left panel of Fig. 3, eigenvalues entangled with irrelevant features are all zero, while eigenvalues which can be traced back to relevant input components are non-zero. Hence, the eigenvalues associated to the input layer can efficiently tag features identified as relevant by the network itself to solve the assigned task. The magnitude of non-zero eigenvalues, i.e., those associated with relevant components, is found to correlate with a direct measure of the relative distance between the distributions from which individual components are drawn (see the right panel of Fig. 3). Since the distributions differ only in their mean value, we quantify the above distance as the absolute value of the difference between the two means. The Pearson correlation ($\rho$) between post-training eigenvalues and the distance between distributions was measured over ten independent training runs, yielding $\rho=0.981\pm 0.008$. This strong correlation provides compelling evidence against the presence of spurious relationships between the two quantities. Given the linear nature of the analyzed dataset, we additionally solved the task using logistic regression, allowing us to compare our results with the statistical measures derived from this model. In particular, we compared the post-training eigenvalues with the coefficients obtained through LASSO regularization, which, in this setup, can serve as an alternative benchmark. We observed a strong agreement between the eigenvalues and the LASSO coefficients, with a very high correlation ($\rho=0.983\pm 0.007$).